# Vehicle Rebalancing Under Adherence Uncertainty

### v1

@@ Abstract
Ride hailing systems suffer from spatial-temporal supply demand imbalance due to drivers operating in independent, and uncoordinated manner. Several fleet rebalancing models have been proposed that can provide repositioning recommendations to idle standing drivers with the objective of maximizing service rate or minimizing customer waiting time. Existing models assume complete adherence by the drivers which leads to limited practical implementation of these models. A novel taxi rebalancing model and a procedure to compute recommendations for repositioning of taxis are proposed which accounts for uncertainties in adherence arising from taxi driver’s preferences and the evolving confidence of the driver in the system due to outcomes of repositioning recommendations. Extensive simulations using NYC taxi dataset showed that the proposed model can lead to 6.3% higher allocation rates, 11.8% higher driver profits, 4.5% higher demand fulfillment, and 14% higher confidence of drivers in the rebalancing system, as compared to a state-of-the-art rebalancing model that is agnostic to the taxi driver preferences and confidence in the system.
@@ I Introduction
Recent urbanization trends have led to increased travel and its associated externalities, such as escalating congestion levels worldwide. In this context, reliance on private cars for personal mobility is becoming increasingly impractical and unsustainable. Ride-hailing services have emerged as an alternative, providing Mobility-on-Demand (MoD) services, but they raise concerns related to the exploitation of public resources, equity, profitability, and, importantly, scalability. Specifically, the travel demand for such services is spatio-temporally asymmetrically distributed (e.g., commuting toward downtown in the morning), making overall operations imbalanced and extremely sensitive to disturbances [1]. Fleet rebalancing has emerged as the solution to address these challenges. Two main classes of approaches have been proposed in the literature [1]. First, researchers have studied mathematical optimization models [2, 3, 4, 5, 6, 7, 8, 9], where upcoming demand is assumed to be known either exactly or approximated using demand prediction models. The optimization model then calculates rebalancing policies based on predicted demands. Second, researchers have employed reinforcement learning techniques to learn taxi rebalancing policies directly from data [10, 11, 12, 13]. Furthermore, from the decision-making perspective, models can be categorized into two classes, i) aggregate models [2, 3, 4, 5], and ii) agent-level models  [6, 7, 8, 9]. The former generate decisions at the fleet-level, determining how many taxis should be moved from one region to another. However, such models do not consider the state of individual taxi drivers, making it challenging to incorporate driver-level objectives such as fairness and preferences. In contrast, agent-level models possess the desired granularity, and offer greater flexibility and control over individual taxis, allowing for tailored movements based on specific criteria or constraints. However, the enhanced flexibility comes at a computational cost, due to the many variables involved in tracking each vehicle. One of the seminal works in the domain of the taxi rebalancing problem proposes an aggregate-level model for a station based taxi fleet management [2]. The objective is to find the rate at which idle taxis should be dispatched from one station to the other to ensure that the queue length of customers waiting at the station remains bounded. Based on the formulation proposed by the authors, the idle-standing taxis are distributed among the stations proportionally to the number of awaiting customers at each station. The excess taxis (if any) are evenly distributed among the stations at a minimum traveling cost. The primary drawback of this work is that it doesn’t consider the upcoming demands while generating the rebalancing decision, and the repositioning decisions are made solely based on the number of customers waiting in the queue. Subsequent work [3] introduced a more complex model, by accounting for both awaiting customers and expected future customer requests in the decision-making process. Taxis are assigned to stations to guarantee that each station has at least as many taxis as there are awaiting customers. Furthermore, following the allocation for awaiting customers, any remaining taxis are distributed proportionally based on demand forecasts at each station. A major shortcoming of [3] is that a simplistic travel demand forecasting assuming demand arrival as a Poisson process. Refer to [14] for a detailed analysis on travel demand forecasting. Secondly, [3] does not account for the uncertainties in the demand forecasts. This is critical, because recommendations might be overly optimistic and cause drivers to head to locations with insufficient demand. This may cause distrust among drivers and hence lead to poor adherence to system recommendations in the future. Authors in [4] further enhanced the aggregate-level taxi rebalancing solution by formulating a multi-period stochastic rebalancing problem considering uncertainties in demand forecasts. The objective of this framework is to ensure that the supply-demand ratio at each station is the same as the overall supply-demand ratio at the network level while minimizing the traveling cost. Formulating the taxi rebalancing problem as a multi-period problem has the advantage of accounting for the downstream effects of decisions made in the planning horizon. As pointed out by the authors, however, this may not always be beneficial due to the underlying uncertainties in the forecasts for mobility patterns. Authors in [5] discuss an upper-confidence-bound based approach to determine optimal parameters, such as the rebalancing frequency, and the length of the planning horizon, to ensure the effective implementation of the aggregate-level models sequentially.
@@
All the models discussed so far are aggregate models and are unable to generate taxi-level decisions. [6] is the agent-level version of the aggregate model introduced by authors in [4], and enables decision-making at the individual taxi level. A notable limitation of this work is its failure to account for uncertainties in travel time and their potential impact on taxi availability at the destination. The authors in [7] greatly improved on this aspect and accounted for the impact of the travel time on taxi availability at the destination station. The availability of a taxi at the destination was represented as a fraction of the planning horizon during which the taxi is expected to be present at the destination, derermined using pre-calculated inter-station travel times. Although [6, 7] present agent-level models, they do not incorporate any driver-centric objectives, rather the objective still remains to balance network-level supply and demand. In our previous work [8], we improved on this aspect by incorporating fairness in providing rebalancing recommendations and [9] and [12] discussed the impact of the charging-related constraints on the taxi rebalancing problem. Apart from vehicle rebalancing. dynamic pricing has also been proposed as a method to mitigate the supply-demand imbalances [15] [16].
@@
Interestingly, research on the impact of human factors on recommender systems is limited. In particular, little attention is given to the analysis of confidence level of taxi drivers and the impact of the rebalancing strategies suggested by the recommender systems on the adherence in the future. Existing taxi rebalancing models are agnostic to taxi driver’s confidence levels as well as their repositioning preferences. Due to this, the rebalancing recommendations given by the model may not be effective in practice because the driver may discard the recommendation if they are not confident about the recommender system or if the recommendation is not aligned with their preference. Driver’s confidence levels are impacted by the outcome of the repositioning recommendation: if the recommendation leads to a higher reward, the confidence level increases and vice-versa. Additionally, at any particular time instant, the confidence level impacts the likelihood of the driver accepting the repositioning recommendation from the system. The dynamics of an agent’s confidence while interacting with a recommender system has been discussed in literature. [17] proposed a Thompson Sampling-based decision-making and confidence update mechanism. This approach has been successfully applied in improving the exploration-exploitation trade-off in recommendation systems [18]. However, the implementation of such methods in taxi recommender systems remains limited. Prior works [19, 20, 21] have also shown that taxi drivers display a preference in terms of their passenger-finding strategies. [21] discussed the driver cruising strategies between the trips and highlights that taxi drivers tend to make reasonable choices between rebalancing and parking, heading to high-demand locations based on the time of day. Authors in [19] used a multinomial logit model to learn the probability distribution over the driver’s choice to head to one of the neighboring zones. Furthermore, authors in [20] presented a model that explores the behavior of taxi drivers when searching for passengers under uncertain conditions and three distinct strategies were studied: random search, maximum anticipated pick-up probability search, and maximum anticipated revenue search. A work that closely relates to our work is by  [22] which discusses a survey based driver preference aware repositioning recommendation model. A notable difference with this work is that it doesn’t consider the impact of the outcome of the repositioning recommendation on the driver’s probability of accepting the recommendation in the future and assumes a static acceptance probability. In our work, we model the driver’s confidence as a dynamic process, which is impacted by repeated interaction with the recommender system.
@@
    We propose to learn the taxi driver preferences using logistic regression and model the taxi driver confidence as a dynamic process using Thompson sampling-based Beta-Bernoulli Bandits. To the best of our knowledge this is the first work to model the taxi drivers’ confidence in the recommender system and its impact on rebalancing recommendations.
@@
    We formulate and solve a novel taxi rebalancing optimization model that predicts the upcoming demands, travel times, driver positions, driver rebalancing preferences, and driver confidence, and uses that as an input to generate taxi rebalancing decisions to maximize profit.
@@
    We analyze the proposed strategy in a simulated taxi network derived from real taxi data from a state-of-the-art New York City dataset, and illustrate its effectiveness in comparison to preference and confidence-agnostic state-of-the-art taxi rebalancing model.
@@ II Mathematical Formulation
The state of each taxi driver is defined as tuple containing i) probability of the driver going to region $j\in\mathcal{R}$ by following his own preference, and ii) probability of the driver accepting the reposition recommendation from the system.
$\left(\mathbb{P}_{c}({\omega_{2}=j|\omega_{1}=0}),\mathbb{P}_{c}({\omega_{1}=1})\right)$ (1)
@@ II-A Taxi Driver’s Preference to Repositioning to a Region
$\mathbb{P}_{c}(\{\omega_{2}=j\}|\{\omega_{1}=0\})=\mathcal{L}_{cj}$ (2)
where $\mathcal{L}_{cj}\in[0,1]$ is defined as follows:
$\mathcal{L}_{cj}=\frac{\sigma_{c}(\mathbf{z_{j}})}{\sum_{i\in\mathcal{R}}\sigma_{c}(\mathbf{z_{i}})}\hskip 7.11317pt\text{and}\hskip 7.11317pt\sigma_{c}(\mathbf{z_{j}})=\frac{1}{1+e^{-\mathbf{w_{c}}\cdot\mathbf{z_{j}}}}$ (3)
Logistic function $\sigma_{c}$ maps a vector of features [^1] The features that have been used in this study, to learn the driver’s repositioning preferences are listed in Fig. 4 of Section IV. $\mathbf{z_{j}}$ associated with region $j\in\mathcal{R}$ to a score which lies in the interval $[0,1]$. This score is indicative of the the attractiveness of region $j\in\mathcal{R}$ for driver $c\in\mathcal{C}$. Parameter $\mathbf{w_{c}}\in\mathbb{R}^{|\mathbf{z_{j}}|+1}$ is unique to each driver $c\in\mathcal{C}$ and is learned by separately fitting a logistic function to the historic repositioning decisions of each driver using the dataset [23].
Alternately, if the driver accepts the system recommendation, i.e., $\omega_{1}=1$, the driver will move to the recommended destination. Once, the driver has accepted the system recommendation, there is no uncertainty in his selection of the destination region.
$\mathbb{P}_{c}({\{\omega_{2}=j\}}|\{\omega_{1}=1\})=x_{cj}$ (4)
where,
@@
Note, $x_{cj}\in\{0,1\}$ is a decision variable, that must be calculated to achieve an optimal dispatch solution. Each taxi driver is recommended only one destination region $j\in\mathcal{R}$:
$\sum_{j\in\mathcal{R}}x_{cj}=1\hskip 2.84544pt\forall c\in\mathcal{C}.$ (6)
@@ II-B Taxi Drivers’ Confidence in the Recommender System
Every driver, $c\in\mathcal{C}$ has an estimate of the probability of success associated with repositioning to region $j\in\mathcal{R}$ when recommended by the system, i.e., $\theta_{r}=\mathbb{P}_{c}(\{\omega_{3}=1\}|\{\omega_{2}=j\},\{\omega_{1}=1\})$, and following his own preference, i.e., $\theta_{p}=\mathbb{P}_{c}(\{\omega_{3}=1\}|\{\omega_{2}=j\},\{\omega_{1}=0\})$. Driver’s estimate of the recommender system’s success probability and following his preference are assumed to be beta distributed and denoted as $\mathbf{\Theta_{r}(c)}\sim\text{Beta}(\alpha_{r}(c),\beta_{r}(c))$ and $\mathbf{\Theta_{p}(c)}\sim\text{Beta}(\alpha_{p}(c),\beta_{p}(c))$ respectively. Beta distribution has a support of $(0,1)$ and the conjugate symmetry property, which makes them suitable for modeling belief distributions [24].
Outcome of the $k^{th}$ recommendation to driver $c\in\mathcal{C}$ is denoted as $y_{c}(k)\in\{0,1\}$ and defined as follows:
@@
$\text{Beta}\left(\alpha_{0}+\epsilon_{1}\sum_{i=1}^{k}y_{c}(i),\hskip 5.69046pt\beta_{0}+\epsilon_{0}\left(k-\sum_{i=1}^{k}y_{c}(i)\right)\right)$
@@
$\mu(c)=\mathbb{P}_{c}(\{\omega_{1}=1\})=\frac{\hat{\theta}_{r}}{\hat{\theta}_{r}+\hat{\theta}_{p}}=\frac{\hat{\theta}_{r}}{M}$ (9)
@@
II-C Driver State Aware Repositioning Recommendation Model
The proposed repositioning recommendation model is used to provide recommendations to the idle-standing taxi drivers. Decision variable $x$ as specified in equation (5) is used to calculate the optimal repositioning recommendations that will maximize the expected profit accumulated by successful allocations of the drivers in the planning horizon.

### v2

@@ Abstract
Ride-hailing systems often suffer from spatiotemporal supply-demand imbalances, largely due to the independent and uncoordinated actions of drivers. While existing fleet rebalancing methods offer repositioning recommendations to idle drivers to improve service efficiency, they typically assume full driver compliance: an unrealistic premise in practice. We propose an Adherence-Aware Vehicle Rebalancing (AAVR) framework that explicitly models and addresses uncertainties in driver adherence, stemming from individual behavioral preferences and dynamic trust in the recommender system. Our approach integrates (i) region-specific XGBoost models for demand forecasting, (ii) a network-level XGBoost model for inter-region travel time prediction, (iii) driver-specific logit models to capture repositioning preferences, and (iv) driver-specific Beta-Bernoulli Bandit models with Thompson Sampling to track and update each driver’s confidence in the system over time. These elements are incorporated into a novel optimization framework that generates adherence-aware repositioning recommendations. To enable real-time implementation, we further develop a linearized version of the AAVR model. Extensive simulations on the NYC taxi dataset demonstrate that AAVR significantly outperforms four state-of-the-art adherence-agnostic baselines, achieving a 28% increase in served demand, a 22.7% reduction in customer wait times, a 28.6% increase in platform earnings and a 26% gain in driver profits on average.
@@ I Introduction
Recent urbanization trends have led to increased travel and its associated externalities, such as increasing levels of congestion worldwide. In this context, the reliance on private cars for personal mobility is becoming increasingly impractical and unsustainable. Ride-hailing services have emerged as an alternative, providing Mobility-on-Demand (MoD) services, but they raise concerns related to the exploitation of public resources, equity, profitability, and, importantly, scalability. Specifically, the travel demand for such services is spatio-temporally asymmetrically distributed (e.g., commuting to the downtown in the morning), making overall operations imbalanced and extremely sensitive to disturbances [1]. Vehicle rebalancing has emerged as a key solution to address these challenges. While several models have been proposed, most assume full compliance from taxi drivers with rebalancing recommendations. In practice, however, drivers may choose to accept or ignore suggestions based on personal preferences or their trust in the system. In this work, we propose an Adherence Aware Vehicle Rebalancing (AAVR) model that explicitly accounts for such behavioral uncertainty. By leveraging historical trip data, AAVR learns driver-specific adherence patterns and incorporates them into the rebalancing optimization process, resulting in more robust and realistic recommendations.
I-A Literature Review
@@
Two main classes of approaches have been proposed in the literature [1]. First, researchers have studied mathematical optimization models [2, 3, 4, 5, 6, 7, 8, 9], where upcoming demand is assumed to be known either exactly or approximated using demand prediction models. The optimization model then calculates rebalancing policies based on predicted demands. Second, researchers have employed reinforcement learning techniques to learn taxi rebalancing policies directly from data [10, 11, 12, 13]. Furthermore, from the decision-making perspective, models can be categorized into two classes, i) aggregate models [2, 3, 4, 5], and ii) agent-level models  [6, 7, 8, 9]. The former generate decisions at the fleet-level, determining how many taxis should be moved from one region to another. However, such models do not consider the state of individual taxi drivers, making it challenging to incorporate driver-level objectives such as fairness and preferences. In contrast, agent-level models possess the desired granularity, and offer greater flexibility and control over individual taxis, allowing for tailored movements based on specific criteria or constraints. However, the enhanced flexibility comes at a computational cost, due to the many variables involved in tracking each vehicle. One of the seminal works in the domain of the taxi rebalancing problem proposes an aggregate-level model for a station based taxi fleet management [2]. The objective is to find the rate at which idle taxis should be dispatched from one station to the other to ensure that the queue length of customers waiting at the station remains bounded. Based on the formulation proposed by the authors, the idle-standing taxis are distributed among the stations proportionally to the number of awaiting customers at each station. The excess taxis (if any) are evenly distributed among the stations at a minimum traveling cost. The primary drawback of this work is that it doesn’t consider the upcoming demands while generating the rebalancing decision, and the repositioning decisions are made solely based on the number of customers waiting in the queue. Subsequent work [3] introduced a more complex model, by accounting for both awaiting customers and expected future customer requests in the decision-making process. Taxis are assigned to stations to guarantee that each station has at least as many taxis as there are awaiting customers. Furthermore, following the allocation for awaiting customers, any remaining taxis are distributed proportionally based on demand forecasts at each station. A major shortcoming of [3] is that a simplistic travel demand forecasting assuming demand arrival as a Poisson process. Secondly, [3] does not account for the uncertainties in the demand forecasts. Refer to [14] for a detailed analysis on travel demand forecasting, and [15] for robust approaches to handle demand uncertainties in the context of the vehicle rebalancing problem. This is critical, because recommendations might be overly optimistic and cause drivers to head to locations with insufficient demand. This may cause distrust among drivers and hence lead to poor adherence to system recommendations in the future. Authors in [4] further enhanced the aggregate-level taxi rebalancing solution by formulating a multi-period stochastic rebalancing problem considering uncertainties in demand forecasts. The objective of this framework is to ensure that the supply-demand ratio at each station is the same as the overall supply-demand ratio at the network level while minimizing the traveling cost. Formulating the taxi rebalancing problem as a multi-period problem has the advantage of accounting for the downstream effects of decisions made in the planning horizon. As pointed out by the authors, however, this may not always be beneficial due to the underlying uncertainties in the forecasts for mobility patterns. Authors in [5] discuss an upper-confidence-bound based approach to determine optimal parameters, such as the rebalancing frequency, and the length of the planning horizon, to ensure the effective implementation of the aggregate-level models sequentially.
@@
All the models discussed so far are aggregate models and are unable to generate taxi-level decisions. [6] is the agent-level version of the aggregate model introduced by authors in [4], and enables decision-making at the individual taxi level. A notable limitation of this work is its failure to account for uncertainties in travel time and their potential impact on taxi availability at the destination. The authors in [7] greatly improved on this aspect and accounted for the impact of the travel time on taxi availability at the destination station. The availability of a taxi at the destination was represented as a fraction of the planning horizon during which the taxi is expected to be present at the destination, determined using pre-calculated inter-station travel times. Although [6, 7] present agent-level models, they do not incorporate any driver-centric objectives, rather the objective still remains to balance network-level supply and demand. In our previous work [8], we improved on this aspect by incorporating fairness in providing rebalancing recommendations and [9, 12] discussed the impact of charging-related constraints on the taxi rebalancing problem. Beyond vehicle rebalancing, dynamic pricing [16] [17], and matching-integrated rebalancing [18], have also been proposed as methods to mitigate imbalances.
Interestingly, research on the impact of human factors on recommender systems is limited. In particular, little attention is given to the analysis of confidence level of taxi drivers and the impact of the rebalancing strategies suggested by the recommender systems on the adherence in the future. Existing taxi rebalancing models are agnostic to taxi driver’s confidence levels as well as their repositioning preferences. Therefore, the rebalancing recommendations given by the model may not be effective in practice because the driver may discard them if they are not confident about the recommender system or if the recommendation is not aligned with their preference. Driver’s confidence levels are impacted by the outcome of the repositioning recommendation: if the recommendation leads to a higher reward, the confidence level increases and vice-versa. Additionally, at any particular time instant, the confidence level impacts the likelihood of the driver accepting the repositioning recommendation from the system.
The evolution of agent’s confidence during interactions with recommender systems has been extensively studied, with growing evidence that human exploratory behavior aligns with Bayesian decision-theoretic models. Building on the foundation of Thompson Sampling [19], researchers have demonstrated significant empirical gains in large-scale recommendation tasks. Controlled experiments further support this view: human behavior in uncertain multi-alternative decision-making tasks is well captured by Thompson Sampling mechanisms [20, 21]. Recent work such as QCARE [22] extends this model by introducing adaptive exploration decay, outperforming utility-based baselines in capturing how humans balance exploration and exploitation. In parallel, collective decision-making has also been interpreted through this lens: Krafft et al.  [23] model group intelligence as a distributed form of Thompson Sampling, where individuals sample actions based on social priors such as popularity, further reinforcing the cognitive plausibility of this mechanism. Despite these advances, Thompson Sampling-based confidence models remain largely unexplored in the context of taxi driver recommender systems. We address this gap by embedding a generalized Beta-Bernoulli Thompson Sampling process into our model, enabling dynamic representation of driver’s evolving confidence in system-generated rebalancing recommendations.
@@
Prior works [24, 25, 26] have also shown that taxi drivers display a preference in terms of their passenger-finding strategies. [26] discussed the driver cruising strategies between the trips and highlights that taxi drivers tend to make reasonable choices between rebalancing and parking, heading to high-demand locations based on the time of day. Authors in [24] used a multinomial logit model to learn the probability distribution over the driver’s choice to head to one of the neighboring zones. Furthermore, authors in [25] presented a model that explores the behavior of taxi drivers when searching for passengers under uncertain conditions and three distinct strategies were studied: random search, maximum anticipated pick-up probability search, and maximum anticipated revenue search. A work that closely relates to our work is by  [27] which discusses a survey based driver preference aware repositioning recommendation model. A notable difference with this work is that it doesn’t consider the impact of the outcome of the repositioning recommendation on the driver’s probability of accepting the recommendation in the future and assumes a static acceptance probability. In our work, we model the driver’s confidence as a dynamic process, which is impacted by repeated interaction with the recommender system.
@@
    We develop a probabilistic framework to model how a taxi driver makes decisions in response to repositioning recommendations. The model captures the sequence of choices a driver makes - whether to accept or reject the recommendation, which region to move to, and whether they are subsequently allocated a passenger. It accounts for differences in individual driver behavior by assigning personalized probabilities to each decision stage.
@@
    Building on the proposed probabilistic framework, we design a learning-based architecture to estimate the key components of each driver’s behavioral model from data. Specifically, we use (i) region-specific XGBoost models to forecast passenger demand; and (ii) a network-level XGBoost model to estimate inter-region travel times; - which are used to estimate allocation probabilities, (iii) driver-specific logit models to estimate the conditional probabilities of region selection when the driver rejects system recommendation; and (iv) Beta-Bernoulli bandit models with Thompson Sampling to infer and update each driver’s probability of accepting recommendations over time, based on their interaction history.
@@
    We formulate an Adherence Aware Vehicle Rebalancing (AAVR) model that integrates driver-specific adherence probabilities into a constrained optimization problem. The original model involves a non-linear objective that jointly maximizes expected allocations and minimizes travel time. We derive an equivalent linearized formulation by introducing auxiliary variables, enabling the problem to be solved efficiently as a mixed-integer linear program (MILP).
  4. 4.
    We analyze the proposed strategy in a simulated taxi network derived from real taxi data from a state-of-the-art New York City dataset, and illustrate its effectiveness in comparison to four adherence-agnostic state-of-the-art taxi rebalancing models and show that these models perform sub-optimally as compared to the proposed model in the uncertain adherence setting.
@@ II Mathematical Formulation
The state of each taxi driver is defined as tuple containing i) probability of the driver going to region $j\in\mathcal{R}$ by following his own preference, and ii) probability of the driver accepting the reposition recommendation from the system:
$\left(\mathbb{P}_{c}({\omega_{2}=j|\omega_{1}=0}),\mathbb{P}_{c}({\omega_{1}=1})\right).$ (1)
@@ II-A Taxi Driver’s Preference to Repositioning to a Region
$\mathbb{P}_{c}(\{\omega_{2}=j\}|\{\omega_{1}=0\})=L_{cj},$ (2)
where $L_{cj}\in[0,1]$ is defined as follows:
$L_{cj}=\frac{\sigma_{c}(\mathbf{z_{j}})}{\sum_{i\in\mathcal{R}}\sigma_{c}(\mathbf{z_{i}})}\hskip 7.11317pt\text{and}\hskip 7.11317pt\sigma_{c}(\mathbf{z_{j}})=\frac{1}{1+e^{-\mathbf{w_{c}}\cdot\mathbf{z_{j}}}}.$ (3)
We use a logit model to learn the repositioning preferences of the taxi drivers. Specifically, a logistic function $\sigma_{c}$ maps a vector of features [^1] The features that have been used in this study, to learn the driver’s repositioning preferences are listed in Fig. 5 of Section IV. $\mathbf{z_{j}}$ associated with region $j\in\mathcal{R}$ to a score which lies in the interval [0,1][0,1][ 0 , 1 ]. This score is indicative of the the attractiveness of region $j\in\mathcal{R}$ for driver $c\in\mathcal{C}$. Parameter $\mathbf{w_{c}}\in\mathbb{R}^{|\mathbf{z_{j}}|+1}$ is unique to each driver $c\in\mathcal{C}$ and is learned by separately fitting a logistic function to the historic repositioning decisions of each driver using the dataset [28].
Alternately, if the driver accepts the system recommendation, i.e., $\omega_{1}=1$, the driver will move to the recommended destination. Once, the driver has accepted the system recommendation, there is no uncertainty in his selection of the destination region:
$\mathbb{P}_{c}({\{\omega_{2}=j\}}|\{\omega_{1}=1\})=\mathbf{x_{cj}},$ (4)
where
@@
Note, $\mathbf{x_{cj}}\in\{0,1\}$ is a decision variable, that must be calculated to achieve an optimal dispatch solution. Each taxi driver is recommended only one destination region $j\in\mathcal{R}$:
$\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}=1\hskip 2.84544pt\forall c\in\mathcal{C}.$ (6)
@@ II-B Taxi Drivers’ Confidence in the Recommender System
Every driver, $c\in\mathcal{C}$ has an estimate of the probability of success associated with repositioning to region $j\in\mathcal{R}$ when recommended by the system, i.e., $\theta_{r}=\mathbb{P}_{c}(\{\omega_{3}=1\}|\{\omega_{2}=j\},\{\omega_{1}=1\})$, and following his own preference, i.e., $\theta_{p}=\mathbb{P}_{c}(\{\omega_{3}=1\}|\{\omega_{2}=j\},\{\omega_{1}=0\})$. Driver’s estimate of the recommender system’s success probability and following his preference are assumed to be beta distributed and denoted as $\mathbf{\Theta_{r}(c)}\sim\text{Beta}(\alpha_{r}(c),\beta_{r}(c))$ and $\mathbf{\Theta_{p}(c)}\sim\text{Beta}(\alpha_{p}(c),\beta_{p}(c))$ respectively. Beta distribution has a support of (0,1)(0,1)( 0 , 1 ) and the conjugate symmetry property, which makes them suitable for modeling belief distributions [29].
The outcome of the $k^{th}$ recommendation to driver $c\in\mathcal{C}$ is denoted as $y_{c}(k)\in\{0,1\}$ and defined as follows:
@@
$\text{Beta}\left(\alpha_{0}+\epsilon_{1}\sum_{i=1}^{k}y_{c}(i),\hskip 5.69046pt\beta_{0}+\epsilon_{0}\left(k-\sum_{i=1}^{k}y_{c}(i)\right)\right).$
@@
$\mu(c)=\mathbb{P}_{c}(\{\omega_{1}=1\})=\frac{\hat{\theta}_{r}}{\hat{\theta}_{r}+\hat{\theta}_{p}}=\frac{\hat{\theta}_{r}}{M}.$ (9)
@@
II-C Adherence Aware Vehicle Rebalancing Model
The proposed model generates repositioning recommendations for idle-standing taxi drivers with the goal of maximizing the total number of successful driver-passenger allocations while minimizing the idle cruising time prior to allocation. It explicitly accounts for uncertainty in driver adherence, influenced by both the driver’s confidence in the recommender system and their individual preference for the suggested destination.